# Discrete Autoregressive Switching Processes in Sparse Graphical Modeling of Multivariate Time Series Data

### v1

@@ 1 Introduction
The rest of the paper is organized as follows. Section 2 introduces the proposed model, including the DAR process and the proposed prior structures, and the MCMC algorithm for posterior inference. Section 3 contains results from the simulation studies and Section 4 illustrates the application to fMRI data on latent learning. Section 5 provides concluding remarks. Julia software is available on GitHub at XXX (to be released upon acceptance).
@@ 2 Sparse Modeling of Multivariate Time Series Data via Cumulative Shrinkage DAR
with state-specific means $\bm{\mu}_{j}$ and precision matrices $\bm{\Omega}_{j}$, $j=1,\ldots,M$, $t=1,\dots,T$. Conditional dependencies can be inferred from the off-diagonal entries of the precision matrices. Specifically, for a given state $j$, if the entry $\omega_{j,il}$ is zero, the corresponding variables $y_{ti}$ and $y_{tl}$ are conditionally independent given the other variables.
@@ 2.1 State Dynamics via Discrete Autoregressive Processes
In order to learn the dependence structure between time points, represented by the sequence $\bm{\gamma}$, we design an approach that employs a discrete autoregressive process, with a cumulative shrinkage prior that enables a computationally efficient estimation of the order of the process. More specifically, we assume that the evolution of the hidden state sequence $\gamma_{t}$ follows a Discrete Autoregressive (DAR) process of order $P$ (Biswas and Song,, 2009), which allows the hidden sequence to incorporate long-term dependencies by considering the previous $P$ lags. Formally, the conditional distribution of $\gamma_{t}$ given the past values $\gamma_{t-1:t-P}$ is expressed as
@@
where $\bm{\phi}=(\phi_{0},\dots,\phi_{P})$ and $\bm{\pi}=(\pi_{1},\dots,\pi_{M})$. We denote with $\{\phi_{j}\}_{j=0}^{P}$ the autoregressive probabilities, with $\phi_{0}=1-\sum_{j=1}^{P}\phi_{j}$, while the state innovation probabilities $\{\pi_{i}\}_{i=1}^{M}$ are defined as $\pi_{i}:=p(\gamma_{t}=i)$, for $i=1,\dots,M$, and allow the process to transition to any of the $M$ states, including those not observed in the previous $P$ lags. Here, $\mathds{1}_{\{j\}}(i)$ is an indicator function equal to 1 if $i=j$ and 0 otherwise. According to (2), the value of $\gamma_{t}$ is chosen as one of the latent states selected at previous time points, $t-1$ to $t-P$, based on the autoregressive probabilities $\{\phi_{j}\}_{j=1}^{P}$. Alternatively, with probability $\phi_{0}$, $\gamma_{t}$ takes on any of the $M$ possible states independently of the history of the latent sequence, according to the state innovation distribution $\bm{\pi}$.
The transition probabilities in the DAR process can be represented by a multi-dimensional array. The dimensions of this array are determined by the number of autoregressive lags, $P$, and the number of hidden states, $M$. As an illustration, when $P=2$, the transition probabilities are described by an $[M\times M\times M]$ array, say $\bm{\eta}$. The individual components of this array, denoted as $\eta_{\,{l,i,j}}$, represent the probability $p(\gamma_{t}=j|\gamma_{t-1}=i,\gamma_{t-2}=l)$ for $l,i,j\in\{1,\dots,M\}$, as defined in (2). However, as the number of lags $P$ increases, the dimensionality of this array grows exponentially. Therefore, the DAR characterization simplifies inference by allowing us to focus only on making inferences on the $\phi$ parameters, as opposed to learning the entire transition matrix, which is the case with HMM models, for example. In fact, when dealing with higher order HMMs, the task involves estimating $M^{P}$ parameters for the transition arrays. In contrast, our proposed method streamlines this process by estimating $(M+P)$ parameters, resulting in a substantial computational advantage.
@@ 2.1.1 Zero-inducing cumulative shrinkage prior for learning time dependence
Proposition 1. Let $\bm{\phi}=\{\phi_{j}\in\Delta_{\phi}:j=0,\dots,P\}$ with $\Delta_{\phi}=\{\phi_{l}:0\leq\phi_{l}\leq 1,\sum_{l=0}^{\infty}=1\}$, be constructed according to (3), and $\bm{v}=\{v_{i}\}_{i=0}^{P}$ and $\bm{z}=\{z_{i}\}_{i=1}^{P}$ be specified as in (4). Under these assumptions, the cumulative shrinkage DAR formulation implies that $p(z_{j+1}=1|z_{j}=1)=1$, for $j=1,\dots,P$.
@@ 2.3 Markov Chain Monte Carlo Algorithm
$p(v_{j}|z_{j})\propto\text{Beta}(a_{v},b_{v})^{z_{j}},\qquad p(z_{j+1}|\bm{v}_{0:j})\propto\text{Bern}(1)^{z_{j}}\,\text{Bern}(\xi_{j})^{1-z_{j}},$ (8)
@@
[…]
@@
    for $t=\hat{P}+1,\dots,T$, and $j_{l}\in\{1,\dots,M\}$, $l=1,\dots\hat{P}$, where $\eta_{\{\,j_{\hat{P}},\dots,j_{1},j_{0}\}}$ are the DAR probabilities of selecting state $j_{0}$, given previous values $j_{1},\dots,j_{\hat{P}}$, as defined in Eq. (2), and $p(\bm{y}_{t}\,|\cdot)$ are the multivariate spiked Gaussian emission densities specified in Eq. (1). Here, we define the backward messages $\beta_{t}(j_{1})=p(\bm{y}_{t:T}|\gamma_{t-1}=j_{1},\cdot)$, as the probability of the partial observation sequence from time $t$ to $T$ given the state $j_{1}$ at time $t-1$, conditioned on all the other parameters. These messages can be recursively expressed as follows (see Proposition 2, Supplementary Material)
@@
We acknowledge that the proposed Bayesian procedure may be susceptible to the label switching problem (Jasra et al.,, 2005) due to the invariance of the likelihood (6) under permutations of the mixture components’ labeling. To mitigate this issue, we adopt a post-processing approach using the Equivalence Classes Representatives (ECR) algorithm, initially introduced by Papastamoulis and Iliopoulos, (2013) and later improved by Rodríguez and Walker, (2014). The core idea of the ECR algorithm is to categorize analogous allocation vectors as mutually exclusive solutions to the label switching problem. In this context, two allocation vectors are considered analogous if one can be obtained from the other merely by permuting its labels. The ECR procedure divides the allocation vectors into analogous categories and identifies a representative from each category. Consequently, during post-processing, the ECR algorithm identifies the permutation corresponding to each MCMC iteration. This permutation is then applied to reorder the matching allocation with the aim of aligning it with the representative of its category.
@@ 2.4 Posterior Inference
where the backward messages are defined as $\beta_{t}(j)=p(\bm{y}_{t:T}|\gamma_{t-1}=j,\cdot)$ and the forward messages are expressed as $\alpha_{t}(j)=p\,(\bm{y}_{1:t-1},\gamma_{t-1}=j\,|\,\cdot)$. In order to leverage the recursive nature of these messages, we also defined the DAR-forward messages $\alpha_{t}(j_{1},\dots,j_{\hat{P}})=p\,(\bm{y}_{1:t-1},\gamma_{t-1}=j_{1},\dots,\gamma_{t-\hat{P}}=j_{\hat{P}}\,|\,\cdot)$. Further details and the validity of these expressions are provided in the Supplementary Material.
@@ 3 Simulation Studies
We investigate the performance of our proposed methodology using simulated data. We wish to assess the ability to recover the true generating DAR probabilities and the emissions parameters, as well as the numbers of autoregressive probabilities and hidden states.
@@ 3.1 Data Generation
(iv) AR(2) graph: in this structure the precision matrix displays an autoregressive pattern of order two over the main diagonal. The entries are specified as $\omega_{il}=\frac{1}{2}$ if $l=i-1,i+1$, $\omega_{il}=\frac{1}{4}$ if $l=i-2,i+2$, and 0 otherwise.
(v) Random sparse graph: for this setting, the precision matrix is generated by randomly selecting $\lfloor\frac{3}{2}D\rfloor$ off-diagonal entries, and drawing each $\omega_{jl}$ uniformly from the interval $[-1.0,-0.4]\cup[0.4,1.0]$, while the diagonal elements are fixed at 1, and the other entries at 0. Each of the off-diagonal element is then divided by the sum of the off-diagonal elements in its row, and then the matrix is averaged with its transpose, to produce a symmetric, positive definite, matrix.
Partial correlation matrices corresponding to these five scenario are displayed in Figure 1 (top row). A single realization from the simulation setting described in this section is shown in Figure 2 (top panel), with vertical colored bands representing the true underlying state sequence. Here, we have further scaled the time series realization, independently for each dimension $d=1,\dots,D$, in such a way that the corresponding standard deviation of those observations $\{y_{td}\}_{t=1}^{T}$ is equal to one. We note that the partial correlations are invariant under a change of scale and origin, allowing a meaningful comparison between true and estimated values of these matrices.
@@ 3.2 Parameter Settings
Figure 2: Simulation Study. (top) time series realization (lines), with each dimension represented by a different colored line; vertical colored bands represent the true underlying state sequence; (bottom) estimated time-varying probability plot.
@@ 3.3 Results
Our approach consistently estimated the correct number of states $\hat{M}$ = 5 as the mode of the posterior distribution and the number of active DAR probabilities as $\hat{P}=2$ with high posterior probability, on all simulated replicates. For a single replicate, in Figure 1 (bottom row) we show the estimates of the state-specific partial correlation matrices, conditioned upon the modal number of states and modal number of DAR parameters. Our approach successfully retrieves the distinct patterns of the true graphs. Figure 2 (bottom panel) displays a time-varying probability plot, namely the local decoding of the hidden state at time $t$, $p(\gamma_{t}=j\,|\,\bm{y},\cdot)$, $j=1,\dots,\hat{M}$, as described in Section 2.4; these plots are constructed by plotting the local probabilities (which add to 1) cumulatively for each $t$, where each state is associated with a different color. The proposed approach appears to correctly retrieve the true latent state sequence. Additionally, Figure 3 displays the posterior histograms of autoregressive and innovation probabilities, with dotted vertical lines denoting the true generating values. Our proposed method appears to provide a good match between true and estimated values for the DAR parameters.
Figure 3: Simulation Study. Posterior histograms of the DAR parameters. (left) autoregressive probabilities $\phi_{l}$, $l=0,\dots\hat{P}$; (right) innovation probabilities $\pi_{j}$, $j=1,\dots,\hat{M}$. Dotted vertical lines denote true parameters. These results are conditioned upon the modal number of states and autoregressive order. The figure is best viewed in colors.

### v2

# Discrete Autoregressive Switching Processes with Cumulative Shrinkage Priors for Graphical Modeling of Time Series Data

@@ 1 Introduction
The rest of the paper is organized as follows. Section 2 introduces the proposed model, including the DAR process and the proposed prior structures, and the MCMC algorithm for posterior inference. Section 3 contains results from the simulation studies and Section 4 illustrates the application to fMRI data on latent learning. Section 5 provides concluding remarks. The Julia software sggmDAR, which implements our proposed methodology is available on GitHub at https://github.com/Beniamino92/sggmDAR.
@@ 2 Sparse Modeling of Multivariate Time Series Data via Cumulative Shrinkage DAR
with state-specific means $\bm{\mu}_{j}$ and precision matrices $\bm{\Omega}_{j}$, $j=1,\ldots,M$, $t=1,\dots,T$. Here, $\mathds{1}_{\{j\}}(\gamma_{t})$ denotes the indicator function, which equals 1 if $\gamma_{t}=j$, and 0 otherwise. Conditional dependencies can be inferred from the off-diagonal entries of the precision matrices. Specifically, for a given state $j$, if the entry $\omega_{j,il}$ is zero, the corresponding variables $y_{ti}$ and $y_{tl}$ are conditionally independent given the other variables.
@@ 2.1 State Dynamics via Discrete Autoregressive Processes
In order to learn the dependence structure between time points, represented by the sequence $\bm{\gamma}$, we design an approach that employs a discrete autoregressive process, with a cumulative shrinkage prior that enables a computationally efficient estimation of the order of the process. More specifically, we assume that the evolution of the hidden state sequence $\gamma_{t}$ follows a Discrete Autoregressive (DAR) process of order $P$ (Biswas and Song,, 2009), a framework originally introduced in the context of multi-lag finite-state Markov chains by Pegram, (1980), and subsequently adapted to categorical time series modeling by Biswas and Song, (2009). The DAR process allows the hidden sequence to incorporate long-term dependencies by considering the previous $P$ lags. Formally, the conditional distribution of $\gamma_{t}$ given the past values $\gamma_{t-1:t-P}$ is expressed as
@@
where $\bm{\phi}=(\phi_{0},\dots,\phi_{P})$ and $\bm{\pi}=(\pi_{1},\dots,\pi_{M})$. We denote with $\{\phi_{j}\}_{j=0}^{P}$ the autoregressive probabilities, with $\phi_{0}=1-\sum_{j=1}^{P}\phi_{j}$, while the state innovation probabilities $\{\pi_{i}\}_{i=1}^{M}$ are defined as $\pi_{i}:=p(\gamma_{t}=i)$, for $i=1,\dots,M$, and allow the process to transition to any of the $M$ states, including those not observed in the previous $P$ lags. Here, $\mathds{1}_{\{j\}}(i)$ is an indicator function equal to 1 if $i=j$ and 0 otherwise. According to Equation (2), at each time point $t$, the model selects a latent state $\gamma_{t}\in\{1,\dots,M\}$, and assigns probabilities based on how this state relates to the past $P$ values of the sequence. Specifically, the probability of choosing a particular state $i\in\{1,\dots,M\}$ is given by the sum of two components: an autoregressive component and an innovation component. The autoregressive probabilities $\{\phi_{j}\}_{j=1}^{P}$ govern the recurrence structure, so that if $\gamma_{t}=\gamma_{t-j}$ for any $j\in\{1,\dots,P\}$, then state $i$ receives mass $\phi_{j}$. That is, the model assigns probability $\phi_{j}$ to each state $i$ that matches $\gamma_{t-j}$. In this way, the latent sequence favors persistence or recurrence of recently visited states. Complementing this, the innovation probability $\phi_{0}$ allows for exploration beyond the observed history. With probability $\phi_{0}$, the model selects a state independently of $\gamma_{t-1},\dots,\gamma_{t-P}$, according to the innovation distribution $\bm{\pi}$, which assigns nonzero probability to each of the $M$ possible states.
The transition probabilities in the DAR process can be represented by a multi-dimensional array. The dimensions of this array are determined by the number of autoregressive lags, $P$, and the number of hidden states, $M$. As an illustration, when $P=2$, the transition probabilities are described by an $M\times M\times M]$ array, say $\bm{\eta}$. The individual components of this array, denoted as $\eta_{\,{l,i,j}}$, represent the probability $p(\gamma_{t}=j|\gamma_{t-1}=i,\gamma_{t-2}=l)$ for $l,i,j\in\{1,\dots,M\}$, as defined in (2). However, as the number of lags $P$ increases, the dimensionality of this array grows exponentially. Therefore, the DAR characterization simplifies inference by allowing us to focus only on making inferences on the $\phi$ parameters, as opposed to learning the entire transition matrix, which is the case with HMM models, for example. In fact, when dealing with higher order HMMs, the task involves estimating $M^{P}$ parameters for the transition arrays. In contrast, our proposed method streamlines this process by estimating $(M+P)$ parameters, resulting in a substantial computational advantage.
@@ 2.1.1 Zero-inducing cumulative shrinkage prior for learning time dependence
Proposition 1. Let $\bm{\phi}=\{\phi_{j}\in\Delta_{\phi}:j=0,\dots,P\}$ with $\Delta_{\phi}=\{\phi_{l}:0\leq\phi_{l}\leq 1,\sum_{l=0}^{\infty}\phi_{l}=1\}$, be constructed according to (3), and $\bm{v}=\{v_{i}\}_{i=0}^{P}$ and $\bm{z}=\{z_{i}\}_{i=1}^{P}$ be specified as in (4). Under these assumptions, the cumulative shrinkage DAR formulation implies that $p(z_{j+1}=1|z_{j}=1)=1$, for $j=1,\dots,P$.
@@ 2.3 Markov Chain Monte Carlo Algorithm
$p(v_{j}|z_{j})\propto\text{Beta}(a_{v},b_{v})^{1-z_{j}},\qquad p(z_{j+1}|\bm{v}_{0:j})\propto\text{Bern}(1)^{z_{j}}\,\text{Bern}(\xi_{j})^{1-z_{j}},$ (8)
@@
  • Update $\bm{\Omega}_{j}$, $\bm{\Lambda}$, and $\bm{\tau}$: We use the augmented block Gibbs sampler method proposed by Li et al., (2019). We center the observations belonging to each state to its current value of the emission mean, $\bm{\mu}_{j}$, and consider a modified set of observations denoted as $\tilde{\bm{Y}}_{j}=\{\bm{y}_{t}-\bm{\mu}_{j}:\gamma_{t}=j\}$. By doing this, we can closely follow the scheme proposed by Li et al., (2019), which assumes zero-mean multivariate normal distributions. We apply the Gibbs sampler independently for each state $j$, fromj=1j=1 to $M_{\text{max}}$ and subsequently update the global shrinkage parameter $\tau_{j}$ and its corresponding augmented parameter $\xi_{j}$. We refer to the reader to Algorithm 1 of Li et al., (2019), for the details of the GHS sampler.
@@
    for $t=\hat{P}+1,\dots,T$, and $j_{l}\in\{1,\dots,M\}$, $l=1,\dots\hat{P}$, where $\eta_{\{\,j_{\hat{P}},\dots,j_{1},j_{0}\}}$ are the DAR probabilities of selecting state $j_{0}$, given previous values $j_{1},\dots,j_{\hat{P}}$, as defined in Eq. (2), and $p(\bm{y}_{t}\,|\cdot)$ are the multivariate spiked Gaussian emission densities specified in Eq. (1). Here, we define the backward messages $\beta_{t}(j_{1})=p(\bm{y}_{t:T}|\gamma_{t-1}=j_{1},\cdot)$, as the probability of the partial observation sequence from time $t$ to $T$ given the state $j_{1}$ at time t1t-1, conditioned on all the other parameters. These messages can be recursively expressed as follows (see Proposition 2, Supplementary Material)
@@
We acknowledge that the proposed Bayesian procedure may be susceptible to the label switching problem (Jasra et al.,, 2005) due to the invariance of the likelihood (6) under permutations of the mixture components’ labeling. To mitigate this issue, we adopt a post-processing approach using the Equivalence Classes Representatives (ECR) algorithm, initially introduced by Papastamoulis and Iliopoulos, (2013) and later improved by Rodr´ıguez and Walker, (2014). The core idea of the ECR algorithm is to categorize analogous allocation vectors as mutually exclusive solutions to the label switching problem. In this context, two allocation vectors are considered analogous if one can be obtained from the other merely by permuting its labels. The ECR procedure divides the allocation vectors into analogous categories and identifies a representative from each category. Consequently, during post-processing, the ECR algorithm identifies the permutation corresponding to each MCMC iteration. This permutation is then applied to reorder the matching allocation with the aim of aligning it with the representative of its category.
@@ 2.4 Posterior Inference
[…]
@@ 3 Simulation Studies
We investigate the performance of our proposed methodology using simulated data, aiming to assess its ability to recover the true DAR probabilities, emission parameters, and the correct number of autoregressive lags and hidden states. We further evaluate the robustness of the approach under model misspecification and its effectiveness on synthetic fMRI data.
@@ 3.1 Data Generation
(iv) AR(2) graph: in this structure the precision matrix displays an autoregressive pattern of order two over the main diagonal. The entries are specified as $\omega_{il}=\frac{1}{2}$ if l=i1,i+1l=i-1,i+1, $\omega_{il}=\frac{1}{4}$ if l=i2,i+2l=i-2,i+2, and 0 otherwise.
(v) Random sparse graph: for this setting, the precision matrix is generated by randomly selecting $\lfloor\frac{3}{2}D\rfloor$ off-diagonal entries, and drawing each $\omega_{jl}$ uniformly from the interval $1.0,-0.4]\cup[0.4,1.0]$, while the diagonal elements are fixed at 1, and the other entries at 0. Each of the off-diagonal element is then divided by the sum of the off-diagonal elements in its row, and then the matrix is averaged with its transpose, to produce a symmetric, positive definite, matrix.
Partial correlation matrices corresponding to these five scenario are displayed in Figure 1 (top row). A single realization from the simulation setting described in this section is shown in Figure 2 (top panel), with vertical bands representing the true underlying state sequence. Here, we have further scaled the time series realization, independently for each dimension $d=1,\dots,D$, in such a way that the corresponding standard deviation of those observations $\{y_{td}\}_{t=1}^{T}$ is equal to one. We note that the partial correlations are invariant under a change of scale and origin, allowing a meaningful comparison between true and estimated values of these matrices.
@@ 3.2 Parameter Settings
Figure 2: Simulation Study. (top) time series realization (lines), with each dimension represented by a different line; vertical bands represent the true underlying state sequence; (bottom) estimated time-varying probability plot.
@@ 3.3 Results
Our approach consistently estimated the correct number of states $\hat{M}$ = 5 as the mode of the posterior distribution and the number of active DAR probabilities as $\hat{P}=2$ with high posterior probability, on all simulated replicates. For a single replicate, in Figure 1 (bottom row) we show the estimates of the state-specific partial correlation matrices, conditioned upon the modal number of states and modal number of DAR parameters. Our approach successfully retrieves the distinct patterns of the true graphs. Figure 2 (bottom panel) displays a time-varying probability plot, namely the local decoding of the hidden state at time $t$, $p(\gamma_{t}=j\,|\,\bm{y},\cdot)$, $j=1,\dots,\hat{M}$, as described in Section 2.4; these plots are constructed by plotting the local probabilities (which add to 1) cumulatively for each $t$. The proposed approach appears to correctly retrieve the true latent state sequence. Additionally, Figure 3 displays the posterior histograms of autoregressive and innovation probabilities, with dotted vertical lines denoting the true generating values. Our proposed method appears to provide a good match between true and estimated values for the DAR parameters. In the Supplementary Material (Section C.1), we further analyze state-specific mean and variance values across different dimensions, revealing distinct patterns in both location and dispersion across states. These findings provide additional support for the model’s ability to capture meaningful heterogeneity in the underlying data distribution.
Figure 3: Simulation Study. Posterior histograms of the DAR parameters. (left) autoregressive probabilities $\phi_{l}$, $l=0,\dots\hat{P}$; (right) innovation probabilities $\pi_{j}$, $j=1,\dots,\hat{M}$. Dotted vertical lines denote true parameters. These results are conditioned upon the modal number of states and autoregressive order.
@@ 3.4 Simulations for Varying $T$ and $D$
3.5 Additional Simulations
We conducted further simulations to assess the robustness of our model under two complementary settings: (i) a model misspecification scenario, where data are generated from a stationary vector autoregressive (VAR) process; and (ii) a realistic synthetic fMRI dataset designed to mimic experimental BOLD signals under structured noise conditions. These simulations provide insight into how the model behaves when its assumptions are violated and how well it performs in capturing task-driven neural dynamics. Full details are provided in the Supplementary Material (Sections F and G).
3.5.1 Misspecified Model
VAR is a popular model for both task and resting state fMRI, commonly used to capture temporal dependencies in neural time series. To explore the model’s robustness under model misspecification, we generated multivariate time series from a stationary first-order VAR process with either sparse or dense autoregressive coefficient matrices. We summarize here the major findings. A complete description of the simulation setup and additional diagnostic plots are provided in the Supplementary Material (Section F).
In the sparse setting, where dependencies among dimensions are limited, the model consistently favored a single latent regime, with an estimated number of states $\hat{M}=1$. This outcome reflects the relative homogeneity of the time series structure under sparsity, which does not induce detectable changes in statistical properties over time. In contrast, under a dense VAR structure, where most components influence one another, the model identified multiple regimes, with $\hat{M}=3$. Despite the absence of true regime-switching in the data-generating process, the inferred latent states captured subtle shifts in the local mean and variance, as well as changes in the estimated connectivity patterns. Residual diagnostics, including standardized residuals and Q-Q plots, confirmed the adequacy of the model fit, with deviations confined to minor departures in the tails. These results highlight the model’s capacity to adapt to a variety of temporal structures, effectively extracting interpretable latent states even under substantial model misspecification.
3.5.2 Simulated fMRI Data
To further assess the performance of our model in a realistic neuroimaging context, we generated synthetic fMRI data using the neuRosim package in R (Welvaert et al.,, 2011), which provides a principled framework for simulating blood-oxygen-level-dependent (BOLD) signals under controlled experimental and noise conditions. This simulation mimics a block-design experiment in which an external stimulus is applied every 80 seconds and persists for 40 seconds. The BOLD response is constructed using a canonical double-gamma hemodynamic response function (HRF), and several structured noise components—namely, temporal drift, physiological oscillations, and scanner-induced artifacts—are independently added to each brain region. The resulting signal is standardized to reflect typical preprocessing steps in fMRI analysis.
Figure 10 displays this process: panel (a) shows the final simulated multivariate fMRI time series; panel (b) presents the clean BOLD signal prior to noise addition; and panel (c) shows the posterior state probability plot obtained from our model. In this setting, the proposed approach correctly infers $\hat{M}=2$ latent states, with state transitions that align closely with the stimulus schedule—despite the unsupervised nature of the analysis. This confirms the model’s ability to recover task-related structure in high-dimensional noisy time series data.
Figure 5: Simulated fMRI Data. (a) Simulated fMRI time series obtained by summing the clean BOLD signal with structured noise components and standardizing each region; (b) True underlying BOLD signal convolved with a canonical double-gamma HRF; (c) Estimated time-varying state probability plot.
The model’s estimated state-specific mean and variance profiles (reported in the Supplementary Material, Section G) highlight interpretable differences between the two inferred states. One state corresponds to higher overall activation across regions, while the other reflects baseline or resting activity. These findings demonstrate the method’s ability to isolate functionally relevant brain states. Additional diagnostics, including standardized residuals and Q-Q plots, further confirm the model’s goodness-of-fit, suggesting that the proposed framework effectively characterizes stimulus-driven neural dynamics even under complex noise conditions.